OpenStudy (anonymous):

Please help,i got 12/0 but IDK if that's right. Find the slope of the line through the pair of points. A(2, –3), P(2, 9)

Parth (parthkohli):

Vertical line, no slope.

OpenStudy (anonymous):

ok here you have two points given A(2, –3), P(2, 9) use the following formula to find slope slope = (y2-y1)/(x2-x1)

OpenStudy (anonymous):

so it would be undefined right? since I got 12/0

OpenStudy (anonymous):

yes you got 12/0 thats right but it results in infinity thus it is a vertical line and the slope for vertical line undefined

OpenStudy (anonymous):

*is

Parth (parthkohli):

Yesh, undefined.

OpenStudy (unklerhaukus):

infinity or negative infinity? @rizwan_uet ,

Parth (parthkohli):

@rizwan_uet \[\dfrac{x}{0} \ne \infty \]

OpenStudy (anonymous):

any number divided by zero = infinity @ParthKohli

OpenStudy (unklerhaukus):

nope

Parth (parthkohli):

Orly?

Parth (parthkohli):

Can you prove it?

OpenStudy (anonymous):

+infinity or -infinity but here both means undefine @UnkleRhaukus

OpenStudy (unklerhaukus):

infinity is defined

Parth (parthkohli):

So you say that it's infinity AND undefined/

OpenStudy (anonymous):

infinity is not defined its not a number you can write. @UnkleRhaukus

Parth (parthkohli):

Infinity is not defined? Are you sure...? Look, I can define it.

OpenStudy (anonymous):

i dont know the proof but i can tell you some thing if you multiply any number by zero it results in zero ok now in your case x/0 . i have to multiply 0 with such a number that results in x so there exists no such number therefore we say that there exist a very large number called infinity such that if we multiply it with zero it will result in x

Parth (parthkohli):

There is no such defined number which exists, thus undefined.

OpenStudy (anonymous):

thats what i said

OpenStudy (anonymous):

undefined is some thing we refer to 0/0 or infinity /infinity thats true but infinity by its self is undefined too. we cant say this point is infinity